OpenStudy (anonymous):

Simplify 6x - 9 - 14x + 8 - 3x. -10x -12x - 9x- 1 -11x- 1

OpenStudy (zepp):

Combine all similar term and make calculations :D

OpenStudy (zepp):

E.g.: in your question, 6x. -14x and -3x are similar term, since they all have an 'x' sticked to them.

OpenStudy (anonymous):

YEA

OpenStudy (zepp):

The important thing here are the signs, you may want to put it like this, so you can understand better and not get confused with the +/-\[(6x) +(- 9)+( - 14x) + (8)+( - 3x)\]

OpenStudy (anonymous):

IS IT D

OpenStudy (anonymous):

its d

OpenStudy (zepp):

Yep :D

OpenStudy (zepp):

\[(6x)+(-14x)+(-3x)+(-9)+(8)=6x-14x-3x-9+8=-8x-3x-1=-11x-1\]

OpenStudy (zepp):

\(-8x-3x-1\\-11x-1\) Sorry